OpenStudy (anonymous):

simplify. (25x²-25)/(1-x)

OpenStudy (anonymous):

take common number 25 from numerator. x^2-1=(x+1)(x-1) so u get 25*(x+1)(x-1)/(1-x)=-25(x+1)

OpenStudy (anonymous):

thank you so much. I was getting stuck at the x^2-1

OpenStudy (anonymous):

;)